### 01 Feb 2024 v1

@@ 5.1 Evaluation of Predictive Performance
Datasets. We conducted a comprehensive set of experiments using a diverse range of real-world time series data obtained from GluonTS (Alexandrov et al., 2020). These datasets are widely recognized for benchmarking time series forecasting models. The determination of the prediction range ($Q$) for each dataset was based on their respective configurations within GluonTS. For each dataset, we performed a sequential split into training, validation, and testing sets. The data was standardized using the mean and standard deviation obtained from each time series within the training set. The predictions were rescaled to their original values when computing evaluation metrics. Please see Appendix §A for dataset details and preprocessing.
@@ 5.2 Model Interpretation
Moreover, we observe the adjustment of correlation strengths facilitated by the dynamic component weights. Specifically, when the weight assigned to the identity matrix ($w_{3}$) increases, the error process tends to exhibit greater independence. On the contrary, when the weights assigned to the other kernel matrices ($w_{0}$, $w_{1}$, and $w_{2}$) are larger, the error process becomes more correlated, since the kernel matrices with different lengthscales are combined to formulate a specific autocorrelation structure. Fig. 3 (a) demonstrates pronounced daily patterns in autocorrelation, particularly when errors exhibit an increased correlation around 17:00 each day. Figs. 3 (b, d) reveal that the autocorrelation at 3:00 in the morning is less pronounced compared to that observed at 17:00. This underscores the crucial need for our methodology to dynamically adapt the covariance matrix, enabling effective modeling of these temporal variations. Figs. 3 (c, e) illustrate the corresponding covariance matrix of the associated target variables within the autocorrelation horizon. The diagonal blocks represent the covariance $\boldsymbol{\Sigma}_{t^{\prime}}$ of time series variables $\mathbf{z}_{t^{\prime}}$ at each time step, which models the correlations between time series. The off-diagonal blocks depict the covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$, $\forall\Delta\neq 0$ within the autocorrelation horizon that can be captured by our approach. The zoomed-in view provides a $3B\times 3B$ region that illustrates the autocorrelation within two lags. We can observe that the autocorrelation is most pronounced at lag 1, which is consistent with the observation of Fig. 3 (a) that the component weight $w_{0}$ assigned to the base kernel matrix with lengthscale $l=1$ is more pronounced than $w_{1}$ and $w_{2}$.
@@ Appendix A Dataset Details
These datasets are widely employed for benchmarking time series forecasting models, following their default configurations in GluonTS, including granularity, prediction range ($Q$) and the number of rolling evaluations. For each dataset, we performed a sequential split into training, validation, and testing sets, with the temporal length of the validation set matching that of the testing set. The temporal length of the testing set was determined by considering the prediction range and the required number of rolling evaluations. For instance, the testing horizon for the $\mathtt{traffic}$ dataset is computed as $24+7-1=30$ time steps. Consequently, the model is tested for predicting 24 steps ($Q$) in a rolling manner with 7 different consecutive prediction start timestamps. In our experiments, we align the conditioning range ($P$) with the prediction range ($Q$), maintaining consistency with the default setting in GluonTS. For simplicity, we set the autocorrelation horizon ($D$) to also match the prediction range ($Q$). Essentially, in this paper, we have $P=Q=D$. The statistics of all datasets are summarized in Table 2.
@@ C.3 Hyperparameters
The batch size for all models is fixed at 32. The LSTM hyperparameters in GPVar are consistent with those outlined in Salinas et al. (2019). This includes utilizing three layers of LSTM, each with 40 nodes, and setting the dropout rate to 0.1. For the distribution head, a single linear layer is employed to output the distribution parameters. A series of distinct linear layers and ELU layers are utilized for generating the component weights. To ensure parity in the number of parameters between Transformer and GPVar, we configure the Transformer with three stacked decoding layers. These layers have a hidden size of 42, $H=2$ attention heads, and a dropout rate of 0.1.

### 07 Feb 2024 v2

@@ 5.1 Evaluation of Predictive Performance
Datasets. We conducted a comprehensive set of experiments using a diverse range of real-world time series data obtained from GluonTS (Alexandrov et al., 2020). These datasets are widely recognized for benchmarking time series forecasting models. The determination of the prediction range ($Q$) for each dataset was based on their respective configurations within GluonTS. We conducted a sequential split into training, validation, and testing sets for each dataset. Each dataset underwent standardization using the mean and standard deviation calculated from the respective time series within the training set. Predictions were then reverted to their original scales for the computation of evaluation metrics. Please see Appendix §A for dataset details and preprocessing.
@@ 5.2 Model Interpretation
Moreover, we observe the adjustment of correlation strengths facilitated by the dynamic component weights. Specifically, as the weight assigned to the identity matrix ($w_{3}$) increases, the error process tends to show greater independence. On the contrary, when the weights assigned to the other kernel matrices ($w_{0}$, $w_{1}$, and $w_{2}$) are larger, the error process becomes more correlated, since the kernel matrices with different lengthscales are combined to formulate a specific autocorrelation structure. Fig. 3 (a) highlights noticeable daily patterns in autocorrelation, especially showcasing heightened correlation among errors around 17:00 each day. Figs. 3 (b, d) indicate that the autocorrelation is less pronounced at 3:00 in the morning compared to what is observed at 17:00. This highlights the important need for our method to adjust the covariance matrix as needed, so we can effectively model the autocorrelation over time. Figs. 3 (c, e) show the covariance matrix of the target variables within the autocorrelation horizon. The diagonal blocks represent the covariance $\boldsymbol{\Sigma}_{t^{\prime}}$ of time series variables $\mathbf{z}_{t^{\prime}}$ at each time step, which models the correlations between time series. The off-diagonal blocks depict the covariance $\operatorname{Cov}(\mathbf{z}_{t-\Delta},\mathbf{z}_{t})$, $\forall\Delta\neq 0$ within the autocorrelation horizon that can be captured by our approach. The zoomed-in view provides a $3B\times 3B$ region that illustrates the autocorrelation within two lags. We can observe that the autocorrelation is most pronounced at lag 1, which is consistent with the observation of Fig. 3 (a) that the component weight $w_{0}$ assigned to the base kernel matrix with lengthscale $l=1$ is more pronounced than $w_{1}$ and $w_{2}$.
@@ Appendix A Dataset Details
These datasets are widely employed for benchmarking time series forecasting models, following their default configurations in GluonTS, including granularity, prediction range ($Q$) and the number of rolling evaluations. We split each dataset sequentially into training, validation, and testing sets. The temporal length of the validation set matched that of the testing set. The length of the testing set was determined based on the prediction range and the number of rolling evaluations required. For instance, the testing horizon for the $\mathtt{traffic}$ dataset is computed as $24+7-1=30$ time steps. Consequently, the model will predict 24 steps ($Q$) in a rolling manner with 7 different consecutive prediction start timestamps. In our experiments, we align the conditioning range ($P$) with the prediction range ($Q$), maintaining consistency with the default setting in GluonTS. For simplicity, we set the autocorrelation horizon ($D$) to also match the prediction range ($Q$). Essentially, in this paper, we have $P=Q=D$. The statistics of all datasets are summarized in Table 2.
@@ C.3 Hyperparameters
The batch size for all models is fixed at 32. The LSTM hyperparameters in GPVar are consistent with those outlined in (Salinas et al., 2019). This includes utilizing three layers of LSTM, each with 40 nodes, and setting the dropout rate to 0.1. For the distribution head, a single linear layer is employed to output the distribution parameters. A series of distinct linear layers and ELU layers are utilized for generating the component weights. To ensure parity in the number of parameters between Transformer and GPVar, we configure the Transformer with three stacked decoding layers. These layers have a hidden size of 42, $H=2$ attention heads, and a dropout rate of 0.1.